# Convex cocompact groups in real hyperbolic spaces with limit set a Pontryagin sphere

### v2

@@ 5. Special subgroups with limit set a Menger curve
Our proof builds on the work in [SW16, DS21, DHW23] and needs the following terminology. We will say that $K$ is inseparable if it is connected and has no separating simplex, no separating pair of non-adjacent vertices, and no separating suspension of a simplex. In other words, $K$ cannot be disconnected by removing any of the following forbidden subcomplexes: a vertex, an edge, a triangle, two non-adjacent vertices, two edges meeting at a vertex, or two triangles meeting along an edge.
@@
Let ${K}$ be a flag-no-square triangulation of a closed connected orientable surface $S$ of genus $g\geqslant 1$. Let $M\subseteq{K}$ be a non-planar inseparable full subcomplex. Let $v\in M^{(0)}$ such that $\operatorname{lk}(v,M)$ is a circle, and let $L$ be the full subcomplex of $M$ spanned by $M^{(0)}\smallsetminus\{v\}$. Then $L$ is a non-planar inseparable full subcomplex with $\operatorname{pcd}(L)=1$.
@@
The statement in (2.) can be deduced from (1.) thanks to [DS21, Theorem 0.1.(2)]. We only need to check that $L$ is not a join. By contradiction, say $L$ is the join of two subcomplexes $L_{1},L_{2}$. Since $L$ has no induced squares and embeds in a surface, we have that (up to reversing the roles) $L_{1}$ is a single vertex and $L_{2}$ embeds in a circle. In particular, it follows that $L$ embeds in a disk, contradicting the fact that $L$ is non-planar. ∎
@@
The complex defining $\Gamma^{6}$ has a collection of $7$ pairwise non-adjacent vertices, given by the $7$ centers of the hexagonal tiling of the defining complex; see Figure 5. By Proposition 5.1, the special subgroup generated by the other $14$ vertices has Gromov boundary homeomorphic to the Menger curve. Since the full subcomplex on these $14$ vertices is $1$-dimensional (i.e., it is a triangle-free graph) the Gromov boundary could also be computed with [DHW23, Corollary 1.7]. Note that this graph has $K_{3,3}$ as a minor, i.e., one can obtain $K_{3,3}$ from this graph by deleting edges, deleting vertices, and contracting edges. So, this graph is non–planar by Wagner’s theorem.
1
@@
The complex defining $\Gamma^{4}$ has a collection of $14$ pairwise non-adjacent vertices, which one can find by trying to fit a hexagonal grid in this complex; see Figure 5. By Proposition 5.1, the special subgroup generated by the other $36$ vertices has Gromov boundary homeomorphic to the Menger curve.
We note that, in contrast to the case of $\Gamma^{6}$, one cannot directly apply [DHW23, Corollary 1.7] to find a special subgroup with limit set a Menger curve, because the complex $K$ defining $\Gamma^{4}$ does not admit a triangle-free inseparable full subgraph. This can be seen as follows. Since every vertex in $K$ has degree $6$, in any triangle-free subgraph $L\subseteq K$, every vertex can have degree at most $3$. On the other hand, if $L$ has a vertex $v$ of degree $1$ or $2$, then the link of $v$ in $L$ is a global cut vertex or a global cut pair respectively. It follows that every vertex of $L$ must have degree exactly $3$. In particular, every edge of $L$ is contained in $2$ edge-loops of length $6$, so the complementary regions of $L$ in $K$ are given by regular hexagons. However, $K$ is not tiled by hexagons, because the number of triangles is $100$, which is not a multiple of $6$.
@@
Bourdon [BO97] previously constructed examples of convex cocompact subgroups of $\mathrm{Isom}(\mathbb{H}^{4})$ with limit set a Menger curve. It was explained to the first-named author by François Guéritaud that at least some of Bourdon’s examples are (perhaps up to deformation) commensurable to hyperbolic reflection groups, as is illustrated by the following example.
@@
Denoting by $s_{1},\ldots,s_{6}$ the corresponding standard generators, consider the index-$6$ reflection subgroup $W^{\prime}$ of $W$ generated by $s_{1}$, $s_{2}$, $t_{1}:=s_{3}$, $t_{2}:=s_{4}$, $t_{3}:=s_{5}s_{3}s_{5}$, $t_{4}:=(s_{5}s_{6}s_{5})s_{4}(s_{5}s_{6}s_{5})$, $t_{5}:=(s_{6}s_{5}s_{6})s_{3}(s_{6}s_{5}s_{6})$, and $t_{6}:=s_{6}s_{4}s_{6}$ (note that the dihedral subgroup of $W$ generated by $s_{5}$ and $s_{6}$ is a retract of $W$, and $W^{\prime}$ is nothing but the kernel of the retraction map from $W$ onto this dihedral subgroup). The subgroup $\Gamma<W^{\prime}$ generated by $s_{1}t_{2k+1}$ and $s_{2}t_{2k+2}$ for $0\leqslant k\leqslant 2$ is of index $4$ in $W^{\prime}$ and is isomorphic to the group $\Gamma_{63}$ in Bourdon’s notation [BO97]; the latter is, as demonstrated by Benakli [BE92], a Gromov hyperbolic group with Gromov boundary a Menger curve. Now the group $W$ can be realized as a convex cocompact reflection group in $\mathrm{Isom}(\mathbb{H}^{4})$; indeed, by considering the signature of the associated Gram matrix, one sees that there is a unique way to do this (up to conjugation) so that the hyperbolic distance between the fixed hyperplane of $s_{3}$ and the fixed hyperplane of $s_{5}$ is equal to the distance between the fixed hyperplane of $s_{4}$ and the fixed hyperplane of $s_{6}$ (and in this case, this distance is $\mathrm{arccosh}(3/\sqrt{2})$). The resulting representation of the subgroup $\Gamma<W$ is of the form described in [BO97].
[…]

### v3

@@ 5. Special subgroups with limit set a Menger curve
Our proof builds on the work in [46, 16, 15] and needs the following terminology. We will say that $K$ is inseparable if it is connected and has no separating simplex, no separating pair of non-adjacent vertices, and no separating suspension of a simplex. In other words, $K$ cannot be disconnected by removing any of the following forbidden subcomplexes: a vertex, an edge, a triangle, two non-adjacent vertices, two edges meeting at a vertex, or two triangles meeting along an edge.
@@
Let K{K} be a flag-no-square triangulation of a closed connected orientable surface $S$ of genus $g\geqslant 1$. Let $M\subseteq{K}$ be a non-planar inseparable full subcomplex. Let $v\in M^{(0)}$ such that $\operatorname{lk}(v,M)$ is a circle, and let $L$ be the full subcomplex of $M$ spanned by $M^{(0)}\smallsetminus\{v\}$. Then $L$ is a non-planar inseparable full subcomplex with $\operatorname{pcd}(L)=1$.
@@
The statement in (2.) can be deduced from (1.) thanks to [16, Theorem 0.1.(2)]. We only need to check that $L$ is not a join. By contradiction, say $L$ is the join of two subcomplexes $L_{1},L_{2}$. Since $L$ has no induced squares and embeds in a surface, we have that (up to reversing the roles) $L_{1}$ is a single vertex and $L_{2}$ embeds in a circle. In particular, it follows that $L$ embeds in a disk, contradicting the fact that $L$ is non-planar. ∎
@@
The complex defining $\Gamma^{6}$ has a collection of $7$ pairwise non-adjacent vertices, given by the $7$ centers of the hexagonal tiling of the defining complex; see Figure 5. By Proposition 5.1, the special subgroup generated by the other $14$ vertices has Gromov boundary homeomorphic to the Menger curve. Since the full subcomplex on these $14$ vertices is $1$-dimensional (i.e., it is a triangle-free graph) the Gromov boundary could also be computed with [15, Corollary 1.7]. Note that this graph has $K_{3,3}$ as a minor, i.e., one can obtain $K_{3,3}$ from this graph by deleting edges, deleting vertices, and contracting edges. So, this graph is non–planar by Wagner’s theorem.
@@
The complex defining $\Gamma^{4}$ has a collection of $15$ pairwise non-adjacent vertices, which one can find by trying to fit a hexagonal grid in this complex; see Figure 5. By Proposition 5.1, the special subgroup generated by the other $35$ vertices has Gromov boundary homeomorphic to the Menger curve.
We note that, in contrast to the case of $\Gamma^{6}$, one cannot directly apply [15, Corollary 1.7] to find a special subgroup with limit set a Menger curve, because the complex $K$ defining $\Gamma^{4}$ does not admit a triangle-free inseparable full subgraph. This can be seen as follows. Since every vertex in $K$ has degree $6$, in any triangle-free subgraph $L\subseteq K$, every vertex can have degree at most $3$. On the other hand, if $L$ has a vertex $v$ of degree $1$ or $2$, then the link of $v$ in $L$ is a global cut vertex or a global cut pair respectively. It follows that every vertex of $L$ must have degree exactly $3$. In particular, every edge of $L$ is contained in $2$ edge-loops of length $6$, so the complementary regions of $L$ in $K$ are given by regular hexagons. However, $K$ is not tiled by hexagons, because the number of triangles is $100$, which is not a multiple of $6$.
@@
Bourdon [8] previously constructed examples of convex cocompact subgroups of $\mathrm{Isom}(\mathbb{H}^{4})$ with limit set a Menger curve. It was explained to the first-named author by François Guéritaud that at least some of Bourdon’s examples are (perhaps up to deformation) commensurable to hyperbolic reflection groups, as is illustrated by the following example.
@@
Denoting by $s_{1},\ldots,s_{6}$ the corresponding standard generators, consider the index-$6$ reflection subgroup $W^{\prime}$ of $W$ generated by $s_{1}$, $s_{2}$, $t_{1}:=s_{3}$, $t_{2}:=s_{4}$, $t_{3}:=s_{5}s_{3}s_{5}$, $t_{4}:=(s_{5}s_{6}s_{5})s_{4}(s_{5}s_{6}s_{5})$, $t_{5}:=(s_{6}s_{5}s_{6})s_{3}(s_{6}s_{5}s_{6})$, and $t_{6}:=s_{6}s_{4}s_{6}$ (note that the dihedral subgroup of $W$ generated by $s_{5}$ and $s_{6}$ is a retract of $W$, and $W^{\prime}$ is nothing but the kernel of the retraction map from $W$ onto this dihedral subgroup). The subgroup $\Gamma<W^{\prime}$ generated by $s_{1}t_{2k+1}$ and $s_{2}t_{2k+2}$ for $0\leqslant k\leqslant 2$ is of index $4$ in $W^{\prime}$ and is isomorphic to the group $\Gamma_{63}$ in Bourdon’s notation [8]; the latter is, as demonstrated by Benakli [4], a Gromov hyperbolic group with Gromov boundary a Menger curve. Now the group $W$ can be realized as a convex cocompact reflection group in $\mathrm{Isom}(\mathbb{H}^{4})$; indeed, by considering the signature of the associated Gram matrix, one sees that there is a unique way to do this (up to conjugation) so that the hyperbolic distance between the fixed hyperplane of $s_{3}$ and the fixed hyperplane of $s_{5}$ is equal to the distance between the fixed hyperplane of $s_{4}$ and the fixed hyperplane of $s_{6}$ (and in this case, this distance is $\mathrm{arccosh}\left(\tfrac{3}{4}\sqrt{2}\right)$. The resulting representation of the subgroup $\Gamma<W$ is of the form described in [8].